2.8) What would be an investor's valuation of a stock with an expected annual dividend in one year of $1.10 if dividends are expected to grow at a constant rate of 10 percent over a long period of time (i.e. forever) and the investor's required return is 15 percent?

a. $20

b. $11

c. $4.40

d) $7.33

e) $22
Answer #1

Solution: 22

Working:

Given: D1 = next year expected annual dividend per share = $1.10;

k = Investor's required rate of return = 0.15;

g = expected dividend growth rate = .10

Value of stock = D1 / (k-g) = 1.10 / (0.15 - 0.10) = 1.10 / 0.05 = 22